1. What is the pH of a solution with [H^ + ] of 1.0*10^ -9 M? What is its pOH? (Kw=[H^ + ][OH^ - ])



2. The [H3O^ + ] of a solution is 1.3 x 10 ^ - 5 , Calculate the [OH^ - ], pH and pOH of the solution. Is the solution acidic or basic?



3. The pH of a solution is 8.7. What are the pOH, [H3O^+] and [OH^-] of the solution?



1
Expert's answer
2022-05-18T05:03:04-0400

1. pH = 9, pOH = 5

2. pH = 4.89, pOH = 9.11, [OH-] = 7.7×10-10

3. pOH = 5.3, [H3O+] = 1.995×10-9, [OH-] = 5.01×10-6